MONT 100N – Modeling the Environment

Please answer the following questions by checking the response that matches your opinion.

1. The pace of the course so far has been:

: too slow : too fast : about right

2. The class lecture notes posted on the course home page have been:

: interesting and useful : not interesting and/or not useful : I have not been reading them

3. The chapter projects so far have been

: interesting and useful

Explain:

: not interesting and/or not useful

Explain:

4. The thing I would most like to see change is: (be specific and suggest what should change and how)